OpenStudy (anonymous):

the algebraic expressions x-2 divided by x² -9is undefined when x is ..?

OpenStudy (anonymous):

It's undefined when the denominator is 0. This occurs when \[x^2-9=0 \rightarrow x=\pm 3\]

OpenStudy (anonymous):

if the denominator is zero than wouldnt the answer be zero?

OpenStudy (anonymous):

No, the denominator is zero for x-values such that \[x^2-9=0\] By what you're saying, if the answer is zero, x=0. But then the denominator would be \[0^2-9=-9\]which is fine. You can divide your expression by -9.

OpenStudy (anonymous):

\[\frac{x-2}{x^2-9}=\frac{something}{0}\]is undefined.

OpenStudy (anonymous):

You have to find the x-values that make the denominator 0.

OpenStudy (anonymous):

Is that clearer?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

Is it the 'undefined' thing, i.e. *why* you have to find x where x^2-9 = 0?

OpenStudy (anonymous):

yea.. i dont understand that..

OpenStudy (anonymous):

Okay. In division, you can't ever divide by 0. This is because, if we did it, we wouldn't actually end up with something that made sense. I'll give you a simple example.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

You're happy with\[\frac{6}{2}=3\] yeah?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So if I then write,\[6=\left[ ? \right]\times 3\]what number would you put in there?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Right. Now,

OpenStudy (anonymous):

if we imagine \[\frac{6}{0}=\left[ ? \right]\](i.e. *some* number)

OpenStudy (anonymous):

we could write

OpenStudy (anonymous):

\[6=0 \times \left[ ? \right]\]

OpenStudy (anonymous):

What number would you put in now?

OpenStudy (anonymous):

thats impossible

OpenStudy (anonymous):

EXACTLY

OpenStudy (anonymous):

We attempted to divide 6 by 0 and ended up with an IMPOSSIBILITY.

OpenStudy (anonymous):

That's why division by zero is undefined.

OpenStudy (anonymous):

Is that better?

OpenStudy (anonymous):

This will happen no matter what you try to divide by zero.

OpenStudy (anonymous):

yeahh

OpenStudy (anonymous):

So that's why you have to find all the x-values that will end up making x^2-9 = 0 ... because everything falls apart there.

OpenStudy (anonymous):

If x is 3 or -3, x^2-9 = 3^2-9 = 9 - 9 = 0 or x^2-9 = (-3)^2-9 = 9-9=0

OpenStudy (anonymous):

\[\frac{x-2}{x^2-9}=\frac{3 -2}{3^2-9}=\frac{1}{0}\]when x=3

OpenStudy (anonymous):

\[\frac{x-2}{x^2-9}=\frac{-3-2}{(-3)^2-9}=\frac{-5}{0}\]when x=-3.

OpenStudy (anonymous):

oooooooooooooh alrighttt tanks

OpenStudy (anonymous):

In both cases, you'll end up with the same rubbish you get with \[\frac{6}{0}\]

OpenStudy (anonymous):

I think it's clicked now :)

OpenStudy (anonymous):

Feel free to become a fan :P

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Good luck with your algebra. Try and link it to numbers if you have trouble.

OpenStudy (anonymous):

okeydokey